# Inferring flavor mixtures in multijet events

### v2

@@ Abstract
Multijet events with heavy-flavors are of central importance at the LHC since many relevant processes—such as $t\bar{t}$, $hh$, $t\bar{t}h$ and others—have a preferred branching ratio for this final state. Current techniques for tackling these processes use hard-assignment selections through $b$-tagging working points, and suffer from systematic uncertainties because of the difficulties in Monte Carlo simulations. We develop a flexible Bayesian mixture model approach to simultaneously infer $b$-tagging score distributions and the flavor mixture composition in the dataset. We model multidimensional jet events, and to enhance estimation efficiency, we design structured priors that leverages the continuity and unimodality of the $b$-tagging score distributions. Remarkably, our method eliminates the need for a parametric assumption and is robust against model misspecification—It works for arbitrarily flexible continuous curves and is better if they are unimodal. We have run a toy inferential process with signal $bbbb$ and backgrounds $bbcc$ and $cccc$, and we find that with a few hundred events we can recover the true mixture fractions of the signal and backgrounds, as well as the true $b$-tagging score distribution curves, despite their arbitrariness and nonparametric shapes. We discuss prospects for taking these findings into a realistic scenario in a physics analysis. The presented results could be a starting point for a different and novel kind of analysis in multijet events, with a scope competitive with current state-of-the-art analyses. We also discuss the possibility of using these results in general cases of signals and backgrounds with approximately known continuous distributions and/or expected unimodality.
@@ 1 Introduction
The Large Hadron Collider (LHC) is one of the greatest machines ever built by mankind. Every knowledge, every available technology, and every to-be-available technology have been fully coordinated and exploited to give rise to a Machine that gives the maximum of it in all of its aspects. This has been confirmed by its outstanding performance along the years, including the discovery of the Higgs boson [1, 2]. Having more than a decade ahead, with a few Standard Model milestones and the chimera of unforeseen discoveries in the horizon, the physics side of the LHC is challenged to exploit to the fullest every bit of data and information available with tools and techniques at the state-of-the-art level. This work is directly aimed at contributing in this direction by making the most of the mutual information contained at the event-by-event level, exploiting prior information in different ways, and also exploring how to reduce systematic uncertainties by novel data-driven methods. The ideas deployed in this article are along the lines of considering maximizing the potential of LHC data as a new frontier in High-Energy Physics, the information frontier.
@@
To this end, we focus on multijet events and the challenge of finding out their composition in terms of classes of possible flavor combinations. This kind of problem occurs often at the LHC and it has the added complication that simulations often have problems reproducing this type of events [3, 4]. We focus on heavy-flavor tagging, which is of main interest at the LHC because of its connection to heavy particles such as the top-quark and the Higgs boson. This kind of problem is relevant for final states such as for instance $t\bar{t}$, $hh$, $tW$ and $Zh$ among others, since the signal and its corresponding backgrounds in the multijet channel end up defining a multijet selection of events with bottom-quark jets, or $b$-jets.
@@
Tagging $b$-jets is a complex area in High Energy Physics (HEP) that seeks to identify jets that have originated in a bottom-quark [5, 6]. It is customary to construct $b$-taggers [7] that receive the features of a given jet in an event and return a ($b$-tagging) score. This score can be considered as being sampled from different probability distribution functions (PDFs) depending on whether the jet has originated in a bottom-quark (b), charm-quark (c), or light-quark/gluon (u, d, s or g). The less overlap there is in these probability distributions, the better the tagger. This kind of problem is usually faced by defining a working-point as a threshold in this score and then performing the difficult and needed task of calibrating on data the true positive rate for $b$-jets and false positive rates for all other non-b jets whose score results are larger than this threshold.
@@
In this work, we explore the possibility of, instead of using calibrated working points, using the available full score PDFs despite knowing that it should be taken as an approximation of a calibrated curve, which is very difficult to achieve. We develop a new statistical estimation procedure by flexible Bayesian modeling [8] and exploiting structured prior information. Simultaneously, this inference process also returns a posterior for the fraction of events of each class present in the dataset, which is of prime interest in any physics analysis.
@@
The rationale for pursuing our goal is as follows. The object of study consists of multijet events, and therefore the dataset consists of tuples of $b$-tagging scores. Since each one of the events consists of some combination of the individual jet flavors (light, charm and bottom), then the likelihood for a given event needs the PDFs for each flavor. Since the true value for these curves is unknown, but we have some prior knowledge of their shape, we set these curves as parameters in the inference problem. In doing these we show how to take advantage of expected features such as continuity and unimodality to be able to infer arbitrary (nonparametric) continuous curves. The classes in the multijet events are each flavor combination allowed by the physical problem, which is usually a few, and in particular much fewer than all the mathematical possible combinations. Henceforth, as a mixture model problem, the likelihood for one event is the sum of the probability for each class times the probability of the given scores in that class. With the likelihood for the dataset one can then infer a density estimation for the parameters, which are the score PDFs for each flavor and the expected fraction of events for each class.
@@
One of the central problems is the inference of the score PDFs for each jet flavor, or individual components, in the mixture model. In this article, we discretize in bins these curves and we study four modeling strategies. The first is to model these PDFs as being sampled from a Dirichlet distribution a priori; the second is as being sampled from a self-normalized Gaussian process; the third as being a weighted mixture of unimodal curves; and the fourth as a strict unimodal using priors from the previous case. The structured priors in these strategies, which come from shape constraints and/or prior regularization, yield an inner structure to the model that leverages the inference results. Statically speaking, these models share the same likelihood while the difference comes from the prior. In contrast to the common misunderstanding of Bayesian inference that either the prior does not matter with enough sample size or the prior needs to be avoided for it comes from subjective specification, here we adopt a hierarchical Bayes approach, where the structure of the prior does help extract more information from a limited amount of data, and we are effectively learning the prior using the observed data via the pattern of the smoothness or the shape. When there is a structured shape of the true score PDFs, it is important for the model to use such structure in the estimation. Standard frequentist statistical tools assume exchangeability between bins, or shrink bin height estimates towards a global model in the absence of data. Our structured prior approach is better able to capture much more of the true score PDF shape especially in bins with limited observations.
@@
This work is divided as follows. In Section 2 we write down the mathematical formulation of the model, its parts, and the aforementioned strategies. We first describe a one-dimensional model and then the $D$-dimensional model. In Section 3 we show the results of the paper. First, we show a toy one-dimensional model that only works, and to some extent, in the unimodal models. Then we show the results for a 4-dimensional case and we infer all relevant features with all four models. Section 4 contains discussions of the results and their prospects, and Section 5 details the conclusions of the work. All the results of this article and the corresponding scripts to obtain them, have been placed for downloading and reproduction in the Github repository [9].
@@ 3.1 Jet flavor tagging
Jet tagging, especially heavy-flavors such as bottom-quark $b$, is an important area in collider physics programs, since its tools and results are used to reconstruct particles such as the top-quark or the Higgs boson. The main objective in this field is to recognize whether an observed QCD jet has originated in a $b$-, $c$-, light-quark, or a gluon. This classification problem has a long history of methods that have been continuously improving over the years. The latest developments consist mainly of a multivariate analysis (usually a neural network) that uses many features in the jet and in the event to assign a score to the jet. Then each kind of jet has a different probability distribution on this score, permitting to filter jets and events in order to perform physics analyses. Creating these tools and algorithms is a sophisticated field driven nowadays mainly by the ATLAS [5] and CMS [6] collaborations at the LHC.
@@
For the sake of simplicity, and for the purposes of the proof-of-concept pursued in this work, we take the distribution curves in Fig. 1 as true distributions from which we sample synthetic data [^3] Observe that in the framework proposed in this article we use as true distributions the ones that would be the (surely) biased priors when addressing the real data.. We sample the data from a probability density function determined by these curves; i.e. we do not use any physics Monte Carlo to sample the data. To emulate a scenario in which one does not have access to the true distributions, as it is in real data, we do not use as input information these true curve values at any point in the analysis, except for assessing the algorithm performance. Depending on the model we use as prior knowledge different combinations of the following features: shifted distribution curves (which do not match the true ones, as it is the case in real scenarios), the continuity and smoothness of these curves, and/or their expected unimodality. To avoid label switching along the inference process we also use that their modes have some sorting in the GN1 ($b$-tagging) score. In the case of multijet events we also use as prior knowledge that the jets in one event cannot be of any flavor combination, but only some specific combinations; as it would be in a real analysis where one knows which are the expected backgrounds and their jets flavor combination. We find that this is fertile prior knowledge since it relates the dependence of the GN1 scores in a given event to the information that one pursues to extract from the dataset.
@@
In order to apply the techniques and models described in Section 2 we use a dataset consisting of $N=500$ events and we divide the GN1 span in 14 bins. We decide to use this number of bins to have a good balance between the smoothness and jaggedness of the histogram for the data [13]. Along this work we indicate each bin with a point in its center and plot lines joining these points. The dataset is created with 80% of its events corresponding to the $c$-jet class and 20% to the $b$-jet class. After binning the data, this one consists of $N=500$ numbers between 1 and 14. See more details in the Github repository [9].
@@
As a second model, we use the Gaussian process model in Section 2.1.2. We use as the mean of the Gaussian the same beta functions as above which are also the dotted lines at the second row (left and central panels) in Fig. 2. The covariance matrix hyperparameters are (along this work) set to $32\,\rho^{2}=$ 3 and $\sigma^{2}=0.25$ for both classes (see Section 2.1.2), and we use an ordered vector [14, 15] in the fifth bin to fix that in this bin the $c$-distribution is greater than the $b$-distribution. We have also tested inferring $\sigma$ and $\rho$ as parameters, but the running time increases considerably with no noticeable improvement in the results, and with some difficulties in the convergence of the chains as indicated by the $\hat{R}$ parameter [16] (see Github [9]). Therefore in this model we are using the prior knowledge of i) shifted distributions with some proximity to their true values and ii) continuity. The results for the inference are plotted in the second row in Fig. 2. We see in the central plot that again there is no noticeable improvement, mainly because of the non-identifiability of the problem. The convergence for the sample fractions has similar results and explanations as the previous case using Dirichlet priors.
@@
Figure 2: Problem of $N$ events with one jet per event, which can be either from class $c$ or $b$. The problem is addressed with the four described models: (from top to bottom) Dirichlet, Gaussian process, Unimodal and Point-Estimate. The left and center columns indicate the priors and posteriors on the $c$- and $b$-distributions, respectively. The right column shows the prior and posterior for the fractions of events with a $c$-jet. As expected from the discussion on the identifiability in the 1D problem, we find practically null improvement in the Dirichlet and Gaussian process models for recovering the posteriors of the $c$- and $b$-distributions. The Unimodal model, starting from an agnostic prior, recovers a posterior with similar shape to the true curves. As we show in the next Section, this performance it is improved with many jets per event since the dependence between them at the event-by-event level provides profitable information for the inferential process.
As a third benchmark model we use the Unimodal model presented in Section 2.1.3. This model has the virtue that exploits i) continuity and ii) (mixture of) unimodality, however it is not guided by a prior mean. In any case, we use prior knowledge to also set an ordered vector as in the Gaussian process case. This helps to avoid label-switching [17]. We see in this model that the posterior approaches the true values in comparison to the priors, even in a one-dimensional problem. This is an important achievement and one of the reasons behind it is that unimodal classes break the non-identifiability if the data is not unimodal.
@@
In the scenario worked out in this Section we have three classes ($cccc$, $ccbb$ and $bbbb$), hence the mixture weights distribution in the sample is conveniently plotted in a two-dimensional simplex (from a three-dimensional space). Therefore to show the true fraction, and the prior and posterior distributions for the fractions in the sample, we work with histograms in the simplex. To define these histograms in the simplex we have divided each class fraction into 11 bins. This yields a total of $11\times 12/2=65$ bins, since the three fractions should add to one. We are using a totally agnostic prior knowledge in the fractions, which assigns to each bin the uniform probability of $1/65\approx 0.015$. We plot colored contour-curves on the simplex based on the counting on these bins. In each simplex plot we indicate how much increases the probability from the prior to the posterior in the true fraction bin.

### v3

@@ Abstract
Multijet events with heavy-flavors are of central importance at the LHC since many relevant processes—such as $t\bar{t}$, $hh$, $t\bar{t}h$ and others—have a preferred branching ratio for this final state. Current techniques for tackling these processes use hard-assignment selections through $b$-tagging working points, and suffer from systematic uncertainties because of the difficulties in Monte Carlo simulations. We develop a flexible Bayesian mixture model approach to simultaneously infer $b$-tagging score distributions and the flavor mixture composition in the dataset. We model multidimensional jet events, and to enhance estimation efficiency, we design structured priors that leverages the continuity and unimodality of the $b$-tagging score distributions. Remarkably, our method eliminates the need for a parametric assumption and is robust against model misspecification—It works for arbitrarily flexible continuous curves and is better if they are unimodal. We have run a toy inferential process with signal $bbbb$ and backgrounds $ccbb$ and $cccc$, and we find that with a few hundred events we can recover the true mixture fractions of the signal and backgrounds, as well as the true $b$-tagging score distribution curves, despite their arbitrariness and nonparametric shapes. We discuss prospects for taking these findings into a realistic scenario in a physics analysis. The presented results could be a starting point for a different and novel kind of analysis in multijet events, with a scope competitive with current state-of-the-art analyses. We also discuss the possibility of using these results in general cases of signals and backgrounds with approximately known continuous distributions and/or expected unimodality.
@@ 1 Introduction
The Large Hadron Collider (LHC) is one of the greatest machines ever built by mankind. State-of-the-art and forward-looking science and technological knowledge were carefully merged to craft a Machine that delivers top performance in all aspects. This has been confirmed by its outstanding performance along the years, including the discovery of the Higgs boson [1, 2]. Having more than a decade ahead, with a few Standard Model milestones and the chimera of unforeseen discoveries in the horizon, the physics side of the LHC is challenged to exploit to the fullest every bit of data and information available with tools and techniques at the state-of-the-art level. This work is directly aimed at contributing in this direction by making the most of the mutual information contained at the event-by-event level, exploiting prior information in different ways, and also exploring how to reduce systematic uncertainties by novel data-driven methods. The ideas deployed in this article are along the lines of considering maximizing the potential of LHC data as a new frontier in High-Energy Physics, the information frontier.
@@
To this end, we focus on multijet events and the challenge of finding their composition in terms of classes of possible flavor combinations. This kind of problem occurs often at the LHC and it has the added complication that simulations often have problems reproducing this type of events [3, 4]. We focus on heavy-flavor tagging ($c$ and $b$ quarks), which is of main interest at the LHC because of its connection to heavy particles such as the top-quark and the Higgs boson. This kind of problem is relevant for final states such as for instance $t\bar{t}$, $hh$, $tW$ and $Zh$ among others, since the signal and its corresponding backgrounds in the multijet channel end up defining a multijet selection of events with bottom-quark jets, or $b$-jets.
@@
Tagging $b$-jets is a complex area in High Energy Physics that seeks to identify jets that have been initiated by a bottom-quark [5, 6]. It is customary to construct $b$-taggers [7] that receive the features of a given jet in an event and return a ($b$-tagging) score. This score can be considered as being sampled from different probability distribution functions (PDFs) depending on whether the jet has been initiated by a bottom-quark (b), charm-quark (c), or light-quark/gluon (u, d, s or g). The less overlap there is in these probability distributions, the better the tagger. This kind of problem is usually faced by defining a working-point as a threshold in this score and then performing the difficult and needed task of calibrating on data the true positive rate for $b$-jets and false positive rates for all other non-$b$ jets whose score results are larger than this threshold.
@@
In this work, we explore the possibility of, instead of using calibrated working points, using the available full score PDFs despite knowing that it should be taken as an approximation of a calibrated curve, which is very difficult to achieve. We develop a new statistical estimation procedure by flexible Bayesian modeling and exploiting structured prior information. Simultaneously, this inference process also returns a posterior for the fraction of events of each class present in the dataset, which is of prime interest in any physics analysis.
@@
The rationale for pursuing our goal is as follows. The object of study consists of multijet events, and therefore the dataset consists of tuples of $b$-tagging scores. Since each one of the events consists of some combination of the individual jet flavors (light, charm and bottom), then the likelihood for a given event needs the PDFs for each flavor. Since the true value for these curves is unknown, but we have some prior knowledge of their shapes, we set these curves as parameters in the inference problem. In doing these we show how to take advantage of physical features such as continuity and unimodality to be able to infer arbitrary (nonparametric) continuous curves. The classes in the multijet events are each flavor combination allowed by the physical problem, which is usually a few, and in particular much fewer than all the mathematical possible combinations. Henceforth, as a mixture model problem, the likelihood for one event is the sum of the probability for each class times the probability of the given scores in that class. With the likelihood for the dataset one can then infer a density estimation for the parameters, which are the score PDFs for each flavor and the fraction of events for each class.
@@
One of the central problems is the inference of the score PDFs for each jet flavor, or individual components, in the mixture model. In this article, we discretize in bins these curves and we study four modeling strategies. The first is to model these PDFs as being sampled from a Dirichlet distribution a priori; the second is as being sampled from a self-normalized Gaussian process; the third as being a weighted mixture of unimodal curves; and the fourth as a strict unimodal using priors from the previous case. The structured priors in these strategies, which come from shape constraints and/or prior regularization, yield an inner structure to the model that leverages the inference results. Statistically speaking, these models share the same likelihood while the difference comes from the prior. In contrast to the common misunderstanding of Bayesian inference that either the prior does not matter with enough sample size or the prior needs to be avoided for it comes from subjective specification, here we adopt a hierarchical Bayes approach, where the structure of the prior does help extract more information from a limited amount of data, and we are effectively learning the prior using the observed data via the pattern of the smoothness or the shape. When there is a structured shape of the true score PDFs, it is important for the model to use such structure in the estimation. Standard frequentist statistical tools assume exchangeability between bins, or shrink bin height estimates towards a global model in the absence of data. Our structured prior approach is better able to capture much more of the true score PDF shape especially in bins with limited observations.
@@
This work is divided as follows. In Section 2 we write down the mathematical formulation of the model, its parts, and the aforementioned strategies. We first describe the one-dimensional model and then the $D$-dimensional model. In Section 3 we show the results of the paper. First, we show a toy one-dimensional model that only works, and to some extent, in the unimodal models. Then we show the results for a four-dimensional case and we infer all relevant features with all four models. Section 4 contains discussions of the results and their prospects, and Section 5 details the conclusions of the work. All the results of this article and the corresponding scripts to obtain them, have been placed for downloading and reproduction in the Github repository [8].
@@ 3.1 Jet flavor tagging
Jet tagging, especially heavy-flavors such as bottom-quark $b$, is an important area in collider physics programs, since its tools and results are used to reconstruct particles such as the top-quark or the Higgs boson. The main objective in this field is to recognize whether an observed QCD jet has been initiated by a $b$-, $c$-, light-quark, or a gluon. This classification problem has a long history of methods that have been continuously improving over the years. The latest developments consist mainly of a multivariate analysis (usually a neural network) that uses many features in the jet and in the event to assign a score to the jet. Then each kind of jet has a different probability distribution on this score, permitting to filter jets and events in order to perform physics analyses. Creating these tools and algorithms is a sophisticated field driven nowadays mainly by the ATLAS [5] and CMS [6] collaborations at the LHC. It should be mentioned that in recent years there has been progress in using pseudo-continuous $b$-tagging (see e.g. [9, 10]), showing the power of using many working points. However, these proposals still work in the same way as a fixed-point calibration, but calibrating in a few (usually $\sim 4$) points.
@@
For the sake of simplicity, and for the purposes of the proof-of-concept pursued in this work, we take the distribution curves in Fig. 1 as true distributions from which we sample synthetic data [^3] Observe that in the framework proposed in this article we use as true distributions the ones that would be the (surely) biased priors when addressing the real data.. We sample the data from a probability density function determined by these curves; i.e. we do not use any physics Monte Carlo to sample the data. To emulate a scenario in which one does not have access to the true distributions, as it is in real data, we do not use as input information these true curve values at any point in the analysis, except for assessing the algorithm performance. Depending on the model we use as prior knowledge different combinations of the following features: shifted distribution curves (which do not match the true ones, as it is the case in real scenarios), the continuity and smoothness of these curves, and/or their expected unimodality. To avoid label switching along the inference process we also use that their modes have some sorting in the GN1 ($b$-tagging) score. (Label switching is a well known problem in mixture models [14] that arises because of a symmetry in the likelihood due to switching of parameters, and is usually addressed by some artificial constraints.) In the case of multijet events we also use as prior knowledge that the jets in one event cannot be of any flavor combination, but only some specific combinations; as it would be in a real analysis where one knows which are the expected backgrounds and their jets flavor combination. We find that this is fertile prior knowledge since it relates the dependence of the GN1 scores in a given event to the information that one pursues to extract from the dataset.
@@
In order to apply the techniques and models described in Section 2 we use a dataset consisting of $N=500$ events and we divide the GN1 span in 14 bins. We decide to use this number of bins to have a good balance between the smoothness and jaggedness of the histogram for the data [15]. Along this work we indicate each bin with a point in its center and plot lines joining these points. The dataset is created with 80% of its events corresponding to the $c$-jet class and 20% to the $b$-jet class. After binning the data, this one consists of $N=500$ numbers between 1 and 14. See more details in the Github repository [8]. We have used the statistical software Stan to perform the inference, and scripts can be found in Ref. [8]; also in Appendix B we describe some relevant details on the hardware and simulation processes.
@@
As a second model, we use the Gaussian process model in Section 2.1.2. We use as the mean of the Gaussian the same beta functions as above which are also the dotted lines at the second row (left and central panels) in Fig. 2. The covariance matrix hyperparameters are (along this work) set to $32\,\rho^{2}=$ 3 and $\sigma^{2}=0.25$ for both classes (see Section 2.1.2), and we use an ordered vector [16, 17] in the fifth bin to fix that in this bin the $c$-distribution is greater than the $b$-distribution. An ordered-vector is a vector-valued random variable that has support on ordered sequences. We have also tested inferring $\sigma$ and $\rho$ as parameters, but the running time increases considerably with no noticeable improvement in the results, and with some difficulties in the convergence of the chains as indicated by the $\hat{R}$ parameter [18] (see Github [8]). Therefore in this model we are using the prior knowledge of i) shifted distributions with some proximity to their true values and ii) continuity. The results for the inference are plotted in the second row in Fig. 2. We see in the central plot that again there is no noticeable improvement, mainly because of the non-identifiability of the problem. The convergence for the sample fractions has similar results and explanations as the previous case using Dirichlet priors.
@@
Figure 2: Problem of $N$ events with one jet per event, which can be either from class $c$ or $b$. The problem is addressed with the four described models: (from top to bottom) Dirichlet, Gaussian process, Unimodal and Point-Estimate. The left and center columns indicate the priors and posteriors on the $c$- and $b$-distributions, respectively. The right column shows the prior and posterior for the fractions of events with a $c$-jet (only posterior is filled). As expected from the discussion on the identifiability in the 1D problem, we find practically null improvement in the Dirichlet and Gaussian process models for recovering the posteriors of the $c$- and $b$-distributions. The Unimodal model, starting from an agnostic prior, recovers a posterior with similar shape to the true curves. As we show in the next Section, this performance it is improved with many jets per event since the dependence between them at the event-by-event level provides profitable information for the inferential process.
As a third benchmark model we use the Unimodal model presented in Section 2.1.3. This model has the virtue that exploits i) continuity and ii) (mixture of) unimodality, however it is not guided by a prior mean. In any case, we use prior knowledge to also set an ordered vector as in the Gaussian process case. This helps to avoid label-switching [14]. We see in this model that the posterior approaches the true values in comparison to the corresponding priors in the left column of the same row, even in a one-dimensional problem. This is an important achievement and one of the reasons behind it is that unimodal classes break the non-identifiability if the data is not unimodal. We find better agreement with the $c$-distribution because it has more data in the sample.
@@
In the scenario worked out in this section we have three classes ($cccc$, $ccbb$ and $bbbb$), hence the mixture weights distribution in the sample is conveniently plotted in a two-dimensional simplex (from a three-dimensional space). Therefore to show the true fraction, and the prior and posterior distributions for the fractions in the sample, we work with histograms in the simplex. To define these histograms in the simplex we have divided each class fraction into 11 bins in the first set of plots (Figs. 36). This yields a total of $11\times 12/2=65$ bins, since the three fractions should add to one. We are using a totally agnostic prior knowledge in the fractions, which assigns to each bin the uniform probability of $1/65\approx 0.015$. We plot colored contour-curves on the simplex based on the counting on these bins. In each simplex plot we indicate how much the density distribution increases when going from the prior to the posterior at the true bin. This is an indicator of how much the posterior is better in estimating the true fraction. In the small signal fraction plots (Fig. 8) we have used 100 bins in order to have a suitable resolution.
@@
Finally, to test the presented framework in more imbalanced conditions and against usual methods, we also run the inference model on samples with 10% and 1% $bbbb$ fractions, and we compare its classification power against cut-based methods.